Adriana is 7 years older than Joshua. The product of their ages is 144. How old is Adriana?

Answer by mananth(16946)

Adriana is 7 years older th yearsan Joshua. The product of their ages is 144. How old is Adriana?
Let Joshua's age be j years
and Adriana's be y years
j+7 =a
j=a-7
j*a = 144
(a-7)*a = 144 ( substitute j)
a^2-7a-144=0
factorises
a^2-16a+9a -144=0
(a-16)(a+9) =0
Takinf the positive value
Adirian's age is 16 years